OpenStudy (coolcat64):

Sasha has a rectangular garden 8 m by 10 m. The garden has a path surrounding it which is 2 m wide. What is the total area of the garden and pathway? Answer Options with 5 options A. 168 m2 B. 120 m2 C. 116 m2 D. 96 m2 E. 84 m2

OpenStudy (coolcat64):

@codingking03

OpenStudy (coolcat64):

@rebeccaxhawaii

OpenStudy (coolcat64):

@southernbelle00

OpenStudy (alanturing):

\[96m^2\]

OpenStudy (codingking03):

helping someone else sorry

OpenStudy (coolcat64):

ok its fine! XD

OpenStudy (alanturing):

The length of the garden is 8 m It's width is 10 m The width of the path is 2 m Total width = 10 + 2 = 12 m Area of rectangle = Length * Width 8 * 12 = 96 Hope it helped you

OpenStudy (coolcat64):

can u help me with a couple of more?

OpenStudy (codingking03):

I am back

OpenStudy (codingking03):

Is all ok?

OpenStudy (codingking03):

or need my help?

OpenStudy (coolcat64):

A diagram shows a rectangular prism with length, 4 centimeters; width, 3 centimeters; and height, 2 centimeters. Question Find the total surface area. Answer Options with 5 options A. 18 cm2 B. 24 cm2 C. 26 cm2 D. 52 cm2 E. none of these

OpenStudy (codingking03):

Id say B

OpenStudy (coolcat64):

The volume of a cube is 216 ft3. What is the length of a side? Answer Options with 5 options A. 6 ft B. 14.7 ft C. 54 ft D. 72 ft E. 108 ft

OpenStudy (codingking03):

hmm

OpenStudy (codingking03):

72

OpenStudy (codingking03):

since its a cube all the sides must be the same...

OpenStudy (codingking03):

and the cube root of 216 is 72 can i have a medal?

OpenStudy (codingking03):

If you can take the other guys medal and give it to me i will give him a medal

OpenStudy (codingking03):

ok?

OpenStudy (coolcat64):

okay

OpenStudy (codingking03):

lol we have the exact same smart score

OpenStudy (coolcat64):

lol